jim and bill bike like to shoot targets . jim ca hit a target with probability of 0.8 , while bill can hit a target with 0.7 . if both fire at a target at the same time , what is the probability that the target is hit at least once?
Answer by jim_thompson5910(35256)

The probability that jim fails to hit the target is 1-0.8 = 0.2


The probability that bill fails to hit the target is 1-0.7 = 0.3


The probability they both fail to hit their target is 0.2*0.3 = 0.06


So the probability that this doesn't happen, ie one of them (or both) succeed, is 1 - 0.06 = 0.94